# Taming the dark photon production via a non-minimal coupling to gravity

### v1

@@
If this bound is violated, $m_{A}$ and $\xi_{1}$ become time-dependent. Therefore, only a small non-minimal coupling is allowed. This result can trivially be generalized to models with multiple scalars.
@@
To summarize, the non-minimal couplings of the Proca field to gravity are meaningful within effective field theory. This theory has a unitarity cutoff that approaches zero in the massless limit, which necessitates model extension at high energies. We have constructed UV completions of the effective theory that are limited by the Planckian physics only. In these completions, the non-minimal couplings are obtained by integrating out heavy spin-0 and spin-2 fields such that $\xi_{i}$ behave as formfactors. The resulting effective theory is limited by the mass scale of these fields, beyond which the formfactors are no longer constant. If one requires the non-minimal couplings to stay constant up to the Hubble scale, their size is constrained to be small. As a result, the theory is ghost- and tachyon-free, and no runaway vector production is allowed. According to the arguments around Eq. 15, this appears to be a general feature of the UV completions.

### v2

@@
If this bound is violated, $m_{A}$ and $\xi_{1}$ become time-dependent. Therefore, only a small non-minimal coupling is allowed. A similar conclusion has been reached in Ref. [14] for the flat space case using dimensional analysis. This result can trivially be generalized to models with multiple scalars.
We note that Ref. [3] has considered a Higgs Abelian extension of the vector non-minimal coupling with $\xi=0$ at tree level. Since gravity preserves gauge symmetry, the loop effects amount to generating gauge invariant corrections to the tree level Lagrangian and a small non-zero $\xi$ is thus induced. Then, our procedure of integrating out a heavy scalar applies and the same conclusion is reached. The difference from the treatment of Ref. [3] is that we keep the external momentum (and the curvature) dependence in the propagators, which is required by the unitarity and EFT considerations. In particular, the non-minimal vector coupling must vanish at large momentum transfer. Our Eq. 7 applies to the special case $\xi|_{\rm tree}=0$ as well.
@@
Eq. 15 can also be used in (nearly) flat space to study interactions of light dark photons. When $H\sim 0$, the EFT description is valid up to the momenta close to the heavy particle threshold $M$. For example, MeV scale dark photons with a substantial non-minimal coupling to gravity represent a viable option at low energies. However, this EFT breaks down in the Early Universe, when the particle momenta become large, and particle production cannot be studied in this approximation.
@@
To summarize, the non-minimal couplings of the Proca field to gravity are meaningful within effective field theory. This theory has a unitarity cutoff that approaches zero in the massless limit, which necessitates model extension at high energies. We have constructed UV completions of the effective theory that are limited by the Planckian physics only. In these completions, the non-minimal couplings are obtained by integrating out heavy spin-0 and spin-2 fields such that $\xi_{i}$ behave as formfactors. The resulting effective theory is limited by the mass scale of these fields, beyond which the formfactors are no longer constant. If one requires the non-minimal couplings to stay constant up to the Hubble scale, their size is constrained to be small. As a result, the theory is ghost- and tachyon-free, and no runaway vector production is allowed.
General unitarity and gauge invariance arguments show that the above features are generic in UV completions and the non-minimal couplings are given by the Ansatz of Eq. 15.